# Bias-Corrected Data Synthesis for Imbalanced Learning

### v1

@@ Appendix B Synthetic Generators
Reweighting is an intuitive oversampling technique used to address imbalanced data in machine learning. This approach works by assigning a higher weight to samples from the minority class so that the training process focuses more on learning from the underrepresented group. For instance, consider a dataset with $n_{1}$ minority samples and $n_{0}$ majority samples with $n_{1}\ll n_{0}$. A common reweighting approach assigns a weight of $w_{1}=\lfloor n_{0}/n_{1}\rfloor$ to each minority sample and a weight of $w_{0}=1$ to each majority sample [Breiman et al., 2017]. This approach is equivalent to oversampling the minority class by replicating each minority sample $\lfloor n_{0}/n_{1}\rfloor-1$ times and training on the resulting augmented dataset with equal weight.
In contrast, bootstrap methods [Efron and Tibshirani, 1994] for imbalanced classification generate synthetic samples by randomly drawing with replacement from the minority samples. Bootstrap can be regarded as a generalization of the fixed-weight reweighting approach as it effectively assigns random weights to the minority samples in each resampling step. While both reweighting and bootstrap are intuitive and straightforward to implement, they are sensitive to outliers in the minority class. By heavily emphasizing or replicating the outliers, these approaches can potentially lead to overfitting to the noise present in the minority group.
@@
Gaussian mixture model [McLachlan and Peel, 2000] is an oversampling technique that assumes the minority samples follow a mixture of multivariate Gaussian distributions with unknown means and covariance matrices. This technique typically fits a single Gaussian component to the minority class. Given the minority samples $\bm{X}_{1},\dots,\bm{X}_{n_{1}}$, this approach first estimates the distributional parameters, including the empirical mean $\hat{\bm{\mu}}_{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\bm{X}_{i}$ and the sample covariance matrix $\hat{\bm{\Sigma}}_{1}=\frac{1}{n_{1}-1}\sum_{i=1}^{n_{1}}(\bm{X}_{i}-\hat{\bm{\mu}}_{1})(\bm{X}_{i}-\hat{\bm{\mu}}_{1})^{\top}$. Next, the synthetic samples are generated by randomly drawing from the estimated Gaussian distribution with mean $\hat{\bm{\mu}}_{1}$ and covariance matrix $\hat{\bm{\Sigma}}_{1}$. This approach effectively captures the first two moments of the minority distribution. However, the strong underlying Gaussian distribution assumption imposes significant constraints. It might generate poorly representative synthetic samples when the true minority distribution is not unimodal, particularly when the distribution is heavy tailed. For example, it involves a non-convex support or is highly skewed. In such cases, the synthetic data fails to accurately reflect the manifold of the minority class. Consequently, the introduction of the synthetic noise can potentially degrade the performance and robustness of the following training step.
@@
SMOTE, introduced by Chawla et al. [2002], is a widely used oversampling method that generates synthetic minority samples in imbalanced datasets. SMOTE generates new synthetic samples by linearly interpolating between pairs of minority samples. It works as follows: for a randomly selected minority class sample, first find its $K$ nearest neighbors in the minority group. Then randomly select one of these $K$ nearest neighbors and create a new point along the line segment between the original point and the chosen neighbor. This procedure is repeated until the desired number of synthetic samples is reached. SMOTE requires a hyperparameter $K$, the number of nearest neighbors considered for each minority sample. Algorithm 2 provides a step-wise description on how SMOTE generates $\tilde{n}_{1}$ synthetic samples based on input data $\bm{X}_{1},\ldots,\bm{X}_{n_{1}}$.
@@
Diffusion models [Ho et al., 2020, Song et al., 2020] form one of the most popular classes of generative models for data synthesis. A diffusion model learns the distribution of observed samples by simulating and statistically revising a Markovian diffusion process that maps the data to standard Gaussian noise and then reconstructs data from noise. The framework consists of two phases: a fixed forward process, which maps a data example to Gaussian noise, and a learned backward process, which iteratively maps random noise back to a realistic data sample.
@@
Flow matching [Lipman et al., 2022] aims to learn a smooth and invertible map from a simple base distribution, say, the standard Gaussian distribution, to the target data distribution. For observation $\bm{x}\in\mathbb{R}^{d}$, consider the probability density path $p:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\int p_{t}(\bm{x}){\rm d}\bm{x}=1$ for any $t\in[0,1]$. Let $p_{0}$ be the simple base distribution, and $p_{1}$ be the target data distribution. Define the flow $\bm{\phi}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ as a time-dependent differomorphic map satisfying that if $\bm{x}\sim p_{0}$, then $\bm{\phi}_{t}(\bm{x})\sim p_{t}$. Without loss of generality, let $\bm{\phi}_{0}(\bm{x})=\bm{x}$. The flow can be generated by a continuous normalizing flow vector field $\bm{v}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that satisfies $\frac{\rm d}{{\rm d}t}\bm{\phi}_{t}(\bm{x})=\bm{v}_{t}(\bm{\phi}_{t}(\bm{x}))$ [Chen et al., 2018]. Flow matching simplifies the learning problem by utilizing conditional flows that define a straight path between the noise and a data point. For a given data point $\bm{x}_{1}\sim p_{1}$ and a noise sample $\bm{x}_{0}\sim p_{0}$, the optimal conditional vector field is the straight line path $\bm{u}_{t}(\bm{x})=\bm{x}_{1}-\bm{x}_{0}$.
@@
$\mathcal{L}_{\rm FM}(\theta)=\mathbb{E}_{t,p_{0},p_{1}}\left[\|v_{t}(\bm{x}_{0}+t(\bm{x}_{1}-\bm{x}_{0});\theta)-(\bm{x}_{1}-\bm{x}_{0})\|^{2}\right],$
@@
There are many other synthetic generators, such as generative adversarial networks (GANs) [Goodfellow et al., 2014, 2020], normalizing flows [Rezende and Mohamed, 2015], and variational autoencoders (VAE) [Kingma and Welling, 2013]. Please see Figueira and Vaz [2022], Lu et al. [2023] for a comprehensive survey.

### v2

@@ Appendix C Synthetic Generators
Reweighting is an intuitive oversampling technique used to address imbalanced data in machine learning. This approach works by assigning a higher weight to samples from the minority class so that the training process focuses more on learning from the underrepresented group. For instance, consider a dataset with $n_{1}$ minority samples and $n_{0}$ majority samples with $n_{1}\ll n_{0}$. A common reweighting approach assigns a weight of $w_{1}=\lfloor n_{0}/n_{1}\rfloor$ to each minority sample and a weight of $w_{0}=1$ to each majority sample (Breiman et al., 2017). This approach is equivalent to oversampling the minority class by replicating each minority sample $\lfloor n_{0}/n_{1}\rfloor-1$ times and training on the resulting augmented dataset with equal weight.
In contrast, bootstrap methods (Efron and Tibshirani, 1994) for imbalanced classification generate synthetic samples by randomly drawing with replacement from the minority samples. Bootstrap can be regarded as a generalization of the fixed-weight reweighting approach as it effectively assigns random weights to the minority samples in each resampling step. While both reweighting and bootstrap are intuitive and straightforward to implement, they are sensitive to outliers in the minority class. By heavily emphasizing or replicating the outliers, these approaches can potentially lead to overfitting to the noise present in the minority group.
@@
Gaussian mixture model (McLachlan and Peel, 2000) is an oversampling technique that assumes the minority samples follow a mixture of multivariate Gaussian distributions with unknown means and covariance matrices. This technique typically fits a single Gaussian component to the minority class. Given the minority samples $\bm{X}_{1},\dots,\bm{X}_{n_{1}}$, this approach first estimates the distributional parameters, including the empirical mean $\widehat{\bm{\mu}}_{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\bm{X}_{i}$ and the sample covariance matrix $\widehat{\bm{\Sigma}}_{1}=\frac{1}{n_{1}-1}\sum_{i=1}^{n_{1}}(\bm{X}_{i}-\widehat{\bm{\mu}}_{1})(\bm{X}_{i}-\widehat{\bm{\mu}}_{1})^{\top}$. Next, the synthetic samples are generated by randomly drawing from the estimated Gaussian distribution with mean $\widehat{\bm{\mu}}_{1}$ and covariance matrix $\widehat{\bm{\Sigma}}_{1}$. This approach effectively captures the first two moments of the minority distribution. However, the strong underlying Gaussian distribution assumption imposes significant constraints. It might generate poorly representative synthetic samples when the true minority distribution is not unimodal, particularly when the distribution is heavy tailed. For example, it involves a non-convex support or is highly skewed. In such cases, the synthetic data fails to accurately reflect the manifold of the minority class. Consequently, the introduction of the synthetic noise can potentially degrade the performance and robustness of the following training step.
@@
SMOTE, introduced by Chawla et al. (2002), is a widely used oversampling method that generates synthetic minority samples in imbalanced datasets. SMOTE generates new synthetic samples by linearly interpolating between pairs of minority samples. It works as follows: for a randomly selected minority class sample, first find its $K$ nearest neighbors in the minority group. Then randomly select one of these $K$ nearest neighbors and create a new point along the line segment between the original point and the chosen neighbor. This procedure is repeated until the desired number of synthetic samples is reached. SMOTE requires a hyperparameter $K$, the number of nearest neighbors considered for each minority sample. Algorithm 2 provides a step-wise description on how SMOTE generates $\tilde{n}_{1}$ synthetic samples based on input data $\bm{X}_{1},\ldots,\bm{X}_{n_{1}}$.
@@
Diffusion models (Ho et al., 2020; Song et al., 2020) form one of the most popular classes of generative models for data synthesis. A diffusion model learns the distribution of observed samples by simulating and statistically revising a Markovian diffusion process that maps the data to standard Gaussian noise and then reconstructs data from noise. The framework consists of two phases: a fixed forward process, which maps a data example to Gaussian noise, and a learned backward process, which iteratively maps random noise back to a realistic data sample.
@@
Flow matching (Lipman et al., 2022) aims to learn a smooth and invertible map from a simple base distribution, say, the standard Gaussian distribution, to the target data distribution. For observation $\bm{x}\in\mathbb{R}^{d}$, consider the probability density path $p:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\int p_{t}(\bm{x}){\rm d}\bm{x}=1$ for any $t\in[0,1]$. Let $p_{0}$ be the simple base distribution, and $p_{1}$ be the target data distribution. Define the flow $\bm{\phi}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ as a time-dependent differomorphic map satisfying that if $\bm{x}\sim p_{0}$, then $\bm{\phi}_{t}(\bm{x})\sim p_{t}$. Without loss of generality, let $\bm{\phi}_{0}(\bm{x})=\bm{x}$. The flow can be generated by a continuous normalizing flow vector field $\bm{v}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that satisfies $\frac{\rm d}{{\rm d}t}\bm{\phi}_{t}(\bm{x})=\bm{v}_{t}(\bm{\phi}_{t}(\bm{x}))$ (Chen et al., 2018). Flow matching simplifies the learning problem by utilizing conditional flows that define a straight path between the noise and a data point. For a given data point $\bm{x}_{1}\sim p_{1}$ and a noise sample $\bm{x}_{0}\sim p_{0}$, the optimal conditional vector field is the straight line path $\bm{u}_{t}(\bm{x})=\bm{x}_{1}-\bm{x}_{0}$.
@@
$L_{\rm FM}(\theta)=\mathbb{E}_{t,p_{0},p_{1}}\left[\|v_{t}(\bm{x}_{0}+t(\bm{x}_{1}-\bm{x}_{0});\theta)-(\bm{x}_{1}-\bm{x}_{0})\|^{2}\right],$
@@
There are many other synthetic generators, such as generative adversarial networks (GANs) (Goodfellow et al., 2014, 2020), normalizing flows (Rezende and Mohamed, 2015), and variational autoencoders (VAE) (Kingma and Welling, 2013). Please see Figueira and Vaz (2022); Lu et al. (2023) for a comprehensive survey.